# Preconditioning without a preconditioner: faster ridge-regression and Gaussian sampling with randomized block Krylov subspace methods

### v1

@@ Acknowledgements
Appendix A Implementation and pseudocode
We describe an implementation of Definition 3.1 based on the block-Lanczos algorithm, which produces a matrix $\mathbf{Q}_{t}$ whose columns form an orthonormal basis for $\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$. A basic version of the block Lanczos algorithm is described in Algorithm 1.
Algorithm 1 Block Lanczos
1:  Input: $\mathbf{A}$, $\mathbf{B}$, $t$
2:  $\hat{\mathbf{Q}}_{0},\hat{\mathbf{B}}_{-1}=\textsc{QR}(\mathbf{B})$, $\hat{\mathbf{Q}}_{-1}=\mathbf{0}$
3:  for $n=0,1,\ldots,t-1$ do
4:     $\mathbf{Y}_{n+1}=\mathbf{A}\hat{\mathbf{Q}}_{n}-\mathbf{Q}_{n-1}\hat{\mathbf{B}}_{n-1}^{\mathsf{T}}$
5:     $\hat{\mathbf{A}}_{n}=\hat{\mathbf{Q}}_{n}^{\mathsf{T}}\mathbf{Y}_{n+1}$
6:     $\mathbf{Z}_{n+1}=\mathbf{Y}_{n+1}-\hat{\mathbf{Q}}_{n}\mathbf{A}_{n}$
7:     for $j=0,1,\ldots,n$ do
8:        $\mathbf{Z}_{n+1}=\mathbf{Z}_{n+1}-\hat{\mathbf{Q}}_{j}(\hat{\mathbf{Q}}_{j}^{\mathsf{T}}\mathbf{Z}_{n+1})$
9:     end for
10:     $\hat{\mathbf{Q}}_{n+1},\hat{\mathbf{B}}_{n}=\textsc{QR}(\mathbf{Z}_{n+1})$
11:  end for
12:  Output: $\{\hat{\mathbf{Q}}_{n}\}$, $\{\hat{\mathbf{A}}_{n}\}$, $\{\hat{\mathbf{B}}_{n}\}$
Note that in line 10 the QR factorization may require deflation if $\mathbf{Z}_{n+1}$ is rank-deficient.
Assuming the algorithm terminates successfully, the output satisfies a symmetric three term recurrence
$\mathbf{A}\hat{\mathbf{Q}}_{n}=\hat{\mathbf{Q}}_{n-1}\hat{\mathbf{B}}_{n-1}^{\mathsf{T}}+\hat{\mathbf{Q}}_{n}\hat{\mathbf{A}}_{n}+\hat{\mathbf{Q}}_{n+1}\hat{\mathbf{B}}_{n}.$ (26)
We obtain the matrix $\mathbf{Q}_{t}$ by assembling the Lanczos basis matrices
$\mathbf{Q}_{t}:=\begin{bmatrix}|&|&&|\\ \hat{\mathbf{Q}}_{0}&\hat{\mathbf{Q}}_{1}&\cdots&\hat{\mathbf{Q}}_{t-1}\\ |&|&&|\\ \end{bmatrix}.$ (27)
Similarly, we obtain a matrix $\mathbf{T}_{t}$ by
$\mathbf{T}_{t}:=\begin{bmatrix}\hat{\mathbf{A}}_{0}&\hat{\mathbf{B}}_{0}^{\mathsf{T}}\\ \hat{\mathbf{B}}_{0}&\hat{\mathbf{A}}_{1}&\ddots&\phantom{\ddots}\\ &\ddots&\ddots&\hat{\mathbf{B}}_{t-2}^{\mathsf{T}}\\ &&\hat{\mathbf{B}}_{t-2}&\hat{\mathbf{A}}_{t-1}\\ \end{bmatrix}.$ (28)
Here the $\hat{\mathbf{B}}_{n}$ are upper triangular, so $\mathbf{T}_{t}$ has total bandwidth $12m+$ 1.
The Krylov basis $\mathbf{Q}_{t}$ and the coefficient matrix $\mathbf{T}_{t}$ are related by Equation 26 which can be written as
$\mathbf{A}\mathbf{Q}_{t}=\mathbf{Q}_{t}\mathbf{T}_{t}+\hat{\mathbf{Q}}_{t}\hat{\mathbf{B}}_{t-1}\mathbf{E}_{t}^{\mathsf{T}},$ (29)
where $\mathbf{E}_{t}$ is the $(mt)\times m$ matrix with the $m\times m$ identity in the last $m$ rows and zeros elsewhere.
The output of the block-Lanczos algorithm can be used to efficiently compute the block-CG iterates.
Theorem A.1.
Let $\mathsf{bcg}_{t}(\mu):=[\mathsf{bcg}_{t}^{(1)}(\mu)~{}\cdots~{}\mathsf{bcg}_{t}^{(m)}(\mu)]$. Then
$\mathsf{bcg}_{t}(\mu)=\mathbf{Q}_{t}(\mathbf{T}_{t}+\mu\mathbf{I})^{-1}\mathbf{E}_{1}\hat{\mathbf{B}}_{-1},$ (30)
where $\mathbf{E}_{1}$ is the $(mt)\times m$ matrix with the $m\times m$ identity in the first $m$ rows and zeros elsewhere.
Proof.
For any matrix $\mathbf{K}$,
$\mathbf{K}(\mathbf{K}^{\mathsf{T}}\mathbf{A}_{\mu}\mathbf{K})^{-1}\mathbf{K}^{\mathsf{T}}\mathbf{b}=\operatornamewithlimits{argmin}_{\mathbf{x}\in\operatorname{range}(\mathbf{K})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$ (31)
Indeed, observe that any $\mathbf{x}\operatorname{range}(\mathbf{K})$ has the form $\mathbf{x}=\mathbf{K}\mathbf{c}$. By direct computation,
$\operatornamewithlimits{argmin}_{\mathbf{c}}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{K}\mathbf{c}\|_{\mathbf{A}_{\mu}}$ (32)
$=\operatornamewithlimits{argmin}_{\mathbf{c}}\|\mathbf{A}_{\mu}^{-1/2}\mathbf{b}-\mathbf{A}_{\mu}^{1/2}\mathbf{K}\mathbf{c}\|_{2}$ (33)
$=((\mathbf{A}_{\mu}^{1/2}\mathbf{K})^{\mathsf{T}}(\mathbf{A}_{\mu}^{1/2}\mathbf{K}))^{-1}(\mathbf{A}_{\mu}^{1/2}\mathbf{K})^{\mathsf{T}}(\mathbf{A}_{\mu}^{-1/2}\mathbf{b})$ (34)
$=(\mathbf{K}^{\mathsf{T}}\mathbf{A}_{\mu}\mathbf{K})^{-1}\mathbf{K}^{\mathsf{T}}\mathbf{b}.$ (35)
This proves Equation 31.
Therefore, by Equation 17, we see that
$\mathsf{bcg}_{t}(\mu)=\mathbf{Q}_{t}(\mathbf{T}_{t}+\mu\mathbf{I})^{-1}\mathbf{Q}_{t}^{\mathsf{T}}\mathbf{B}.$ (36)
We get the final result by noting that $\mathbf{B}=\hat{\mathbf{Q}}_{0}\hat{\mathbf{B}}_{-1}$ and $\mathbf{Q}_{t}^{\mathsf{T}}\hat{\mathbf{Q}}_{0}=\mathbf{E}_{1}$. ∎
The special case in Equation 18 is a consequence of the fact that $\hat{\mathbf{B}}_{-1}$ is obtained as the “R” factor of the QR factorization of $\mathbf{B}$. In particular, the first column is zero except the top entry of $\hat{\mathbf{B}}_{-1}$ which is $\|\mathbf{b}\|$.
Proof of Theorem 3.7.
At each iteration the block-Lanczos algorithm (Algorithm 1) performs one matrix-load ($m$ parallel matrix products) and lower-order arithmetic dominated by the reorthogonalization costs, in which $\mathbf{Z}_{n+1}$ is orthgonalized against all of the previous $\hat{\mathbf{Q}}_{j}$. Each product $\hat{\mathbf{Q}}_{j}(\hat{\mathbf{Q}}_{j}^{\mathsf{T}}\mathbf{Z}_{n+1})$ costs $O(dm^{2})$ operations, and there are $n$ total products. This is done for iterations $n=0,\ldots,t-1$, resulting in a total arithmetic cost $O(dm^{2}t^{2})$.
Since $\mathbf{T}_{t}$ (and hence $\mathbf{T}_{t}+\mu\mathbf{I}$) are $(mt)\times(mt)$ matrices of bandwidth $O(m)$, the linear system $(\mathbf{T}_{t}+\mu\mathbf{I})\mathbf{z}=\mathbf{e}_{1}$ can be solved in $O((mt)m^{2})=O(m^{3}t)$ time. Subsequently, since $\mathbf{Q}_{t}$ is a $d\times(mt)$ matrix, $\mathbf{Q}_{t}\mathbf{z}$ can be computed in $O(dmt)$ time. ∎
In exact arithmetic, it is in fact possible to reduce the floating point cost to $O(dmtL+m^{3}tL)$ and the storage to $O(dm+dL)$, where $L$ is the number of values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ (O’Leary, 1980). However, there are several caveats to such an implementation. First, such methods require knowing the values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ ahead of time. Second, the finite-precision behavior of such methods can differ greatly from the exact arithmetic behavior. We discuss the impacts of finite precision arithmetic in Section A.1, and in Appendix E we perform some numerical experiments that indicate that reorthogonalization may be unnecessary in some situations.
A.1 Finite precision arithmetic
It is well-known that the behavior of KSMs in finite precision arithmetic may be very different from exact arithmetic (Greenbaum, 1997; Meurant, 2006).
In exact arithmetic, lines 7 and 8 of Algorithm 1 are unnecessary as $\mathbf{Z}_{n+1}$ will already be orthogonal to the existing Krylov subspace. Omitting these lines reduces the arithmetic costs, and also allows the previous columns of the block Krylov subspace to be processed and discarded as they are generated, thereby reducing storage costs as well. In finite precision arithmetic, omitting reorthogonalization can have a major impact. In particular, the columns of $\mathbf{Q}_{t}$ can lose orthogonality (and even linear independence).
For methods based on the standard Lanczos algorithm (Algorithm 1 with block size one), there is quite a bit of theory which guarantees that such methods can still work, even without orthogonalization (Paige, 1971, 1976; Druskin & Knizhnerman, 1991; Greenbaum, 1989; Musco et al., 2018; Chen, 2024).
Unfortunately, much less theory is known about the block-Lanczos algorithm in finite precision arithmetic. In fact, even with reorthogonalization, the behavior of the block-Lanczos algorithm is more much more complicated than the standard Lanczos algorithm. For instance, the QR factorization in line 10 must do some sort of deflation if $\mathbf{Z}_{n+1}$ is (numerically) rank-deficient.
Without reorthogonalization, we observe that in many cases the block-CG method fails to converge at all, while in other cases it does converge for some time; see Appendix E. Understanding this behavior is well beyond the scope of the current work.
Appendix B Omitted proofs
@@
B.1 Nestendness
We begin by proving that the Krylov subspace used by PCG, for certain types of preconditioners, is contained in a certain block-Krylov subspace.
Proof of Theorem 4.1.
$\mathcal{K}_{1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})$ $=\operatorname{span}(\mathbf{P}^{-1}\mathbf{b})$ (37)
$=\operatorname{span}((\mathbf{I}+\mathbf{X})\mathbf{b})$ (38)
$=\operatorname{span}(\mathbf{b}+\mathbf{X}\mathbf{b}).$ (39)
Now, assume that
$\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})$ $\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{(t-1)+s}(\mathbf{A},\mathbf{\Omega}).$
Now we consider the order $t$ subspace
$\hskip 50.00008pt(\mathbf{P}^{-1}\mathbf{A})^{t-2}\mathbf{P}^{-1}\mathbf{b},(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}\}$ (40)
From the inductive hypothesis, we know that for all $j=0,...,t-2$,
$(\mathbf{P}^{-1}\mathbf{A})^{j}\mathbf{P}^{-1}\mathbf{b}$ $\in\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})$ (41)
$\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$ (42)
$\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}),$ (43)
$(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}$ $=(\mathbf{P}^{-1}\mathbf{A})\mathbf{v}$ (44)
$=(\mathbf{I}+\mathbf{X})\mathbf{A}\mathbf{v}$ (45)
$=\mathbf{A}\mathbf{v}+\mathbf{X}\mathbf{A}\mathbf{v}.$ (46)
Clearly $\mathbf{X}\mathbf{A}\mathbf{v}\subset\operatorname{range}(\mathbf{X})\subseteq\mathbf{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. Moreover, as noted above, $\mathbf{v}\in\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$ and so
$\mathbf{A}\mathbf{v}$ $\in\mathbf{A}\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathbf{A}\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$ (47)
$=\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}).$ (48)
This proves the result. ∎
B.2 Nystrom PCG
Proof of Theorem 4.3.
@@ B.2 Nystrom PCG
$\mathbf{E}=\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle,\quad\widehat{\mathbf{A}}_{\mu}=\mathbf{A}\langle\mathbf{K}\rangle+\mu\mathbf{I}.$ (49)
@@
$\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$ (50)
@@
$\widehat{\mathbf{A}}_{\mu}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}+\mu\mathbf{I}=\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$ (51)
@@
$\|\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|$
$\leq\|\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|+\|\mathbf{P}_{\mu}^{-1/2}\mathbf{E}\mathbf{P}_{\mu}^{-1/2}\|.$ (52)
@@
$\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}=(\theta+\mu)\mathbf{U}\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$ (53)
@@
$\|\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|\leq\theta+\mu+\|\mathbf{E}\|.$ (54)
@@
$\|(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})^{-1}\|$ (55)
$=\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|$ (56)
$=\Big{\|}\mathbf{A}_{\mu}^{-1/2}\Big{(}\frac{1}{\theta+\mu}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}$
$\hskip 70.0001pt+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\Big{)}\mathbf{A}_{\mu}^{-1/2}\Big{\|}$ (57)
$\leq\frac{1}{\theta+\mu}\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|$
$\hskip 50.00008pt+\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|.$ (58)
@@
$\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|\leq 1.$ (59)
Likewise, since $\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}\preceq 1$,
$\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|\leq\|\mathbf{A}_{\mu}^{-1}\|=\frac{1}{\lambda_{n}+\mu}.$ (60)
$\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|\leq\frac{1}{\theta+\mu}+\frac{1}{\lambda_{n}+\mu}.$ (61)

### v2

@@ Acknowledgements
Appendix A More numerical experiments
A.1 Convergence
We provide more test problems comparing the convergence of block-CG with Nyström PCG and CG as described in Section 7.1. As observed in Fig. 2, block-CG outperforms the other methods in terms of matrix-loads, and often significantly so.
Note that on the outliers20 problem we now append $\ell=10$ or $\ell=22$ Gaussian vectors (rather than $\ell=20$ as shown in Fig. 2). This problem has 20 large eigenvalues, and the small amount of oversampling significantly improves the convergence of Nyström PCG with $s=1$. We also include a test with the bottom20 problem, which has 20 small eigenvalues.
Figure 5: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on several test problems; see also Fig. 2.
A.2 Block size
Theorem 4.3 suggests the accuracy of low-rank approximations to $\mathbf{A}$ built from the information in the block Krylov subspace $\mathcal{K}_{t}(\mathbf{A},\mathbf{\Omega}_{b})$ depends on $bk$, regardless of the individual value of $b$ and $k$. In Fig. 6, we show the error (after a fixed number of matrix-loads) as a function of the block size, and in Fig. 7 we show the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$ as a function of the block size. In some cases Nyström PCG with depth $s=3$ behaves similarly to Nyström PCG with depth $s=1$ when the same number of matrix-vector products are used. For instance, on the outliers20 problem, both methods have a significant drop in the error (due to a significant drop in the condition number) when dimension of the Krylov subspace is roughly equal to $20$, the number of outlying eigenvalues.
Figure 6: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus the Nyström starting block size $\ell$ (number of columns in $\mathbf{\Omega}$) for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on several test problems.
A.3 Matrix-vector cost
In Fig. 8 we show convergence as a function of matrix-vector products on the same test problems as in Fig. 2. Here our block-CG underperforms the other methods. This does not conflict with any of our theory, which is in terms of matrix-loads. However, it serves as a reminder that our method may not be suitable in all computational settings; see Section 1.2.
Figure 8: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-products for block-CG (Refer to caption), CG (Refer to caption), and Nyström PCG with $s=1$ (Refer to caption) and $s=3$ (Refer to caption) on the same test problems as Fig. 2.
A.4 Test problems
In Fig. 9 we show the spectrums of the test problems we used.
Figure 9: Spectrum of test problems used in this paper.
Appendix B Further discussion on Nyström PCG/CG/BCG
B.1 Bounds for CG in terms of the deflated condition number
In Theorem 4.7 we show that if $s=O(\log(d))$, then $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})=O(\lambda_{r+1}/\lambda_{d})$. It is informative to see bounds for CG in terms of $\lambda_{r+1}/\lambda_{d}$ as well as in terms of the effective dimension. Towards this end, we recall a standard bound for CG; see e.g. [29].
Theorem B.1.
For any $r\geq 0$ the CG iterate satisfies
$\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{cg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{2(t-r)}{\sqrt{\kappa_{r+1}(\mu)}}\right).$
Theorem B.1 guarantees that CG converges at a rate of $\sqrt{\lambda_{r+1}/\lambda_{d}}$ after a burn-in period of $r$ iterations. This is reminiscent of the bounds for Nyström PCG presented in this paper, which guarantee convergence at this rate after a preconditioner build time of roughly $O(r)$ matrix-products.
B.2 Bounds in terms of the effective dimension
The main theoretical bounds for Nyström PCG from [23] are in terms of the effective dimension
(51) $d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl{(}\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr{)}=\sum_{i=1}^{d}\frac{\lambda_{i}}{\lambda_{i}+\mu}.$
Theorem B.2 (Theorem 1.1 in [23]).
(52) $\operatorname{\mathbb{E}}\Bigl{[}\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\Bigr{]}<28.$
We can convert Theorem B.1 to a bound in terms of the effective dimension by noting that all but the first $O(d_{\textup{eff}})$ eigenvalues of $\mathbf{A}$ are bounded by $\mu$. In particular, [23, Lemma 5.4] implies
(53) $r>2d_{\textup{eff}}(\mu)\quad\Longrightarrow\quad\lambda_{r+1}\leq\mu.$
Together, Theorems B.1 and 53 give a bound for CG in terms of the effective dimension.
Corollary B.3.
The CG iterate satisfies
$\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{cg}_{t}\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{cg}_{0}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\sqrt{2}(t-2d_{\textup{eff}}(\mu))\right).$
Corollary B.3 asserts that after a burn in of $2d_{\textup{eff}}(\mu)$ iterations, CG converges in $O(\log(1/\varepsilon))$ iterations. Thus, the theory in [23] does not guarantee Nyström PCG has any major benefits over CG in terms of matrix-vector products.
B.3 Is Nyström PCG a good idea?
The existence of a bound like Theorem B.1 seems damning; CG automatically satisfies bounds similar to those of of Nyström PCG, without the need to construct and store a preconditioner. However, as we now discuss, the full story is much more subtle, and we believe Nyström PCG is still a viable method in many instances.
First, Nyström PCG is able to parallelize matrix-vector products used to build the preconditioner. Thus, the number of matrix-loads can be considerably less than required by CG. This is reflected in our experiments in Fig. 2.
However, these same plots indicated that in terms of matrix-products, CG significantly outperforms Nyström PCG (the $s\ell$ matrix-products to build the Nyström preconditioner require only one matrix-load). On the other hand, in apparent contradiction to our experiments, the experiments in [23] indicate that Nyström PCG consistently and significantly outperforms CG.
Recall that the experiments in Section 7 are done using full-reorthogonalization (so that exact arithmetic theory is still applicable). On the other hand, the experiments in [23] appear to have been done without reorthogonalzation. It’s well-known that Theorem B.1 does not hold finite precision arithmetic without reorthogonalization for any r>0r>0.666A weaker version does still hold [28]. On the other hand, Corollary 2.3 can be expected to hold to close degree [28, 42]. Since Nyström PCG works by explicitly decreasing the condition number of $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$, the convergence of Nyström PCG in finite precision arithmetic is still well described by the exact arithmetic theory for Nyström PCG.
(a) CG (Refer to caption) and Nyström PCG with $s=1$ (Refer to caption) without any reorthgonalization.
(b) CG (Refer to caption) with reorthogonalization for $22$ iterations and Nyström PCG with $s=1$ (Refer to caption) without any reorthgonalization.
Figure 10: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ as a function of matrix products. Light curves show convergence with full-reorthgonalization.
We illustrate this point of comparison explicitly in Fig. 10(a). We run CG and Nyström PCG on the fastdecay problem (for which $\lambda_{1}/\lambda_{d}\gg\lambda_{20}/\lambda_{d}$). As expected, CG with reorthogonalization burns in for roughly 20 iterations, and then begins converging much more rapidly at a rate depending on $\lambda_{21}/\lambda_{d}$, outperforming Nyström PCG with reorthogonalization ($s=1,\ell=22$), which uses 22 matrix-vector products [^7] We need some oversampling to capture the top eigenspace. to build the preconditioner, at which point also converges at a rate depending on $\lambda_{21}/\lambda_{d}$. However, if reorthogonalization is not used, CG converges at a much slower rate depending on $\lambda_{1}/\lambda_{d}$ while Nyström PCG still converges at the faster rate depending on $\lambda_{21}/\lambda_{d}$.
B.4 What if CG uses some reorthogonalization?
Comparing vanilla CG with Nyström PCG as in Fig. 10(a) is perhaps a bit unfair to CG; Nyström PCG incurs costs in building, storing, and applying the preconditioner. To put the algorithms on more equal footing, we might allow CG to use a similar amount of storage and arithmetic operations to do some form of reorthogonalization. While there are many complicated schemes based on detailed analyses of the Lanczos algorithm in finite precision arithmetic [51, 58], a simple scheme is to perform re-orthogonlization only for the first $s\ell$ iterations, and then to continue orthogonalize against these vectors in subsequent iterations. In this case the cost of reorthogonalization during the first $s\ell$ iterations will be $O(ds^{2}\ell^{2})$ operations (the same as the cost to build the the Nyström preconditioner) and subsequently orthogonalizing against these vectors will require $O(ds\ell)$ operations per iteration, the same as the cost to apply the Nyström preconditioner to a vector.
In Fig. 10(b) we show the behavior of CG with full-reorthogonalization for the first 30 iterations and then orthogonalizing subsequent Krylov basis vectors against these first thirty vectors. Compared to Fig. 10(a), we observe that CG now converges as rapidly as Nyström PCG.
B.5 block-CG
In Fig. 2 we compare the performance of these methods when we do not use full-reorthogonalzation. In particular, we run Nyström PCG with $s=3$ without any reorthogonalization. Building the preconditioner requires orthogonalizing $s\ell$ vectors. For block-CG and CG use full reorthogonalization until we have a set of $3\ell$ vectors and then continue to orthogonalize only against these vectors. For reference, we also display the convergence of block-CG and CG if no reorthogonalization is used at all. Unfortunately, far less is known about the behavior of block-CG and block-Lanczos in finite precision arithmetic than is known about their single-vector counterparts.
Figure 11: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ versus matrix-loads for block-CG (Refer to caption) with reorthogonalization for 3 iterations, CG (Refer to caption) with reorthogonalization for $3\ell$ iterations, and Nyström PCG with $s=3$ (Refer to caption) without any reorthgonalization. Light curves show convergence of block-CG and CG with no reorthgonalization. The test problems are the same as Fig. 2.
Appendix C Omitted proofs
@@
C.1 Proof of Proposition 4.1
@@
(54) $\mathbf{E}=\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle,\quad\widehat{\mathbf{A}}_{\mu}=\mathbf{A}\langle\mathbf{K}\rangle+\mu\mathbf{I}.$
@@
(55) $\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
@@
(56) $\widehat{\mathbf{A}}_{\mu}=\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}+\mu\mathbf{I}=\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$
@@
(57) $\|\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|\leq\|\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|+\|\mathbf{P}_{\mu}^{-1/2}\mathbf{E}\mathbf{P}_{\mu}^{-1/2}\|.$
@@
(58) $\mathbf{P}_{\mu}^{-1/2}\widehat{\mathbf{A}}_{\mu}\mathbf{P}_{\mu}^{-1/2}=(\theta+\mu)\mathbf{U}\mathbf{U}^{\mathsf{T}}+\mu(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
@@
(59) $\|\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}\|\leq\theta+\mu+\|\mathbf{E}\|.$
@@
(60) $\hskip-30.00005pt\|(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})^{-1}\|$
(61) $=\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|$
(62) $=\Big{\|}\mathbf{A}_{\mu}^{-1/2}\Big{(}\frac{1}{\theta+\mu}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\Big{)}\mathbf{A}_{\mu}^{-1/2}\Big{\|}$
(63) $\leq\frac{1}{\theta+\mu}\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|+\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|.$
@@
(64) $\|\mathbf{A}_{\mu}^{-1/2}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}\mathbf{A}_{\mu}^{-1/2}\|\leq 1.$
Likewise, since $\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}\preceq\mathbf{I}$,
(65) $\|\mathbf{A}_{\mu}^{-1/2}(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}})\mathbf{A}_{\mu}^{-1/2}\|\leq\|\mathbf{A}_{\mu}^{-1}\|=\frac{1}{\lambda_{n}+\mu}.$
(66) $\|\mathbf{A}_{\mu}^{-1/2}\mathbf{P}_{\mu}\mathbf{A}_{\mu}^{-1/2}\|\leq\frac{1}{\theta+\mu}+\frac{1}{\lambda_{n}+\mu}.$